A 10-card hand is dealt from an ordinary deck of 52 cards. Find the

probability that there are exactly 4 cards of one suit and 3 in two other

suits.
Answer #1

probability that there are exactly 4 cards of one suit and 3 in two other suits

=P(select a suit from 4 and then select 4 cards from it and then select 2 suits from remaining 3 and chooses 3 cards from each of the suits) =4C1*13C4*3C2*13C3*13C3/52C10 =4*715*3*286*286/15820024220 =0.0443621